# Detection of Electron Paramagnetic Resonance of Two Electron Spins Using a Single NV Center in Diamond

### v1

@@ IV Summary
In summary, we demonstrated the detection of EPR from two electron spins using a single NV center in diamond. CPMG-DEER and DEER-Rabi sequences are used to study the spin interactions. The observed NV-EPR signal is in absence of strong hyperfine peaks, and further measurement of Rabi oscillations of the signal suggests the interaction is of a single NV center with two electron spins. The physical system and method presented in this paper allows easy identification of single defect center coupled to a small number of spins and characterization of the interaction strength. The analysis of NV-EPR signal presented can be a method to distinguish the number of interacting spins up to five, which is limited by the detection time window of the oscillations.
The non-equal but measurable interaction strength is favorable in studying non-equilibrium quantum dynamics. Investigation of quantum phase dynamics, like spin liquids or discrete time crystalline (DTC), requires interacting spin system with disordered structures Broholm et al. (2020); Bussandri et al. (2024). Cu2+ ions formed kagome lattice Norman (2016) and Ru3+ ions formed honeycomb lattice Takagi et al. (2019) has been actively researched materials for realization of spin liquid. This multi electron spin system could be a material to look for these emergent quantum effects with the similarity of being a spin 1/2 magnetic disordered type of material. A chain of 9 13C nuclear spins in diamond has been shown to be a programmable spin-based DTC Randall et al. (2021). The characterizable electron spin system in this study is potentially a smaller sized system to realize DTC. A good knowledge of the interaction strengths in a spin system will also be significant in better characterization of non-classical correlations, namely quantum entanglement, which can be key in modeling open quantum dynamics and developing entangled quantum sensing. Quantum enhanced NMR with a 9.4-fold of sensitivity increase has been demonstrated by entangling 9 proton spins with a 31P spin Jones et al. (2009). The electron spins detected in this study can also be simultaneously controlled and potentially achieve entanglement enhancement beyond the standard quantum limit. In addition, further study can be done to reveal more information about the nature of the spin species and improve the understanding of decoherence of shallow NV centers, paving the pathway to higher standard engineering for quantum sensing applications.
@@ Acknowledgements.
This work was supported by the National Science Foundation (ECCS-2204667 and CHE-2004252 with partial co-funding from the Quantum Information Science program in the Division of Physics), the USC Anton B. Burg Foundation, and the Searle scholars program (ST).Data Availability

### v2

@@ IV Summary
In summary, we demonstrated the detection of EPR from two electron spins using a single NV center in diamond. CPMG-DEER and DEER-Rabi sequences are used to study the spin interactions. The observed NV-EPR signal is in the absence of strong hyperfine peaks, and further measurement of Rabi oscillations of the signal suggests the interaction is of a single NV center with two electron spins. The physical system and method presented in this paper allow easy identification of a single defect center coupled to a small number of spins and characterization of the interaction strength. The analysis of the NV-EPR signal presented can be a method to distinguish the number of interacting spins up to five, which is limited by the detection time window of the oscillations.
The non-equal but measurable interaction strength is favorable in studying non-equilibrium quantum dynamics. Investigation of quantum phase dynamics, like spin liquids or discrete time crystalline (DTC), requires interacting spin systems with disordered structures.Broholm et al. (2020); Bussandri et al. (2024) Cu2+ ions formed kagome latticeNorman (2016) and Ru3+ ions formed honeycomb latticeTakagi et al. (2019) has been actively researched materials for realization of spin liquid. This multi electron spin system could be a material to look for these emergent quantum effects with the similarity of being a spin 1/2 magnetic disordered type of material. A chain of 9 13C nuclear spins in diamond has been shown to be a programmable spin-based DTC.Randall et al. (2021) The characterizable electron spin system in this study is potentially a smaller-sized system to realize DTC. A good knowledge of the interaction strengths in a spin system will also be significant in better characterization of non-classical correlations, namely quantum entanglement, which can be key in modeling open quantum dynamics and developing entangled quantum sensing. Quantum-enhanced NMR with a 9.4-fold sensitivity increase has been demonstrated by entangling 9 proton spins with a 31P spin.Jones et al. (2009) The electron spins detected in this study can also be simultaneously controlled and potentially achieve entanglement enhancement beyond the standard quantum limit. In addition, further study can be done to reveal more information about the nature of the spin species and improve the understanding of the decoherence of shallow NV centers, paving the pathway to higher standard engineering for quantum sensing applications.
V Supplementary Material
The supplementary material includes a discussion and simulation of the ESEEM effect observed in the experiment and additional measurement results on the other NV centers.
@@ Acknowledgements.
This work was supported by the National Science Foundation (ECCS-2204667, CHE-2404463 and CHE-2004252 with partial co-funding from the Quantum Information Science program in the Division of Physics), the USC Dornsife, the USC Anton B. Burg Foundation, and the Searle scholars program (ST).Data Availability